### v1

@@ 3.1.2 Domain-specific enhancement based on fine-tuning
      where $\Delta W$ is the Kronecker product of two low-rank matrices $L_{1}$ and $L_{2}$:
      $\Delta W=L_{1}\otimes L_{2}\text{.}$ (15)
      In this formula, $L_{1}$ and $L_{2}$ are two low-rank matrices, and they compute a large matrix that is updated during the fine-tuning process through Kronecker product. LoKr is particularly suitable for tasks that require increasing the model’s dimensions to capture more complex relationships, while maintaining similar parameter efficiency to LoHa. However, LoKr may require more complex mathematical operations to handle the Kronecker product, and in some cases, its computational cost may be higher than that of LoHa.
@@
    Full fine-tuning: Full fine-tuning is not constrained by pre-training tasks or data distributions, making it flexible to adapt to various downstream tasks. It allows the model to be directly optimized end-to-end on the data of the final task without the need for additional adapter modules. However, because it requires updating all parameters in the model, full fine-tuning demands significant computational resources and longer training times. Moreover, foundation models have a vast number of parameters, and insufficient fine-tuning data may lead to overfitting. Additionally, the intermediate variables generated during full fine-tuning consume a large amount of GPU memory. Therefore, researchers have proposed many parameter-efficient fine-tuning methods mentioned earlier, which can reduce resource consumption and training time while maintaining performance [77].
@@ 3.2 Customization of the foundation model based on pre-trained modules
Foundation models may contain millions or even billions of parameters, and through transfer learning, it’s possible to reduce parts of the model that need training, thereby significantly lowering training costs. This approach is known as the customization of the foundation model based on pre-trained modules. The essence of transfer learning lies in utilizing the knowledge embedded in model parameters during the pre-training process when constructing a new model. As mentioned earlier, in the architecture of foundation models, there are typically five main modules: modality encoders, backbone calculators, modality decoders, input projectors, and output projectors. Among them, the modality encoder, backbone calculator, and modality decoder carry a large amount of knowledge as they are directly involved in the encoding, processing, and decoding of data. In contrast, the input projector and output projector themselves carry less model knowledge. In some cases, they may not even have explicit models responsible for these functions, or these modules are only trained when building new foundation models. Therefore, when customizing foundation models, we generally do not choose to transfer the input projector and output projector.
@@
Next, this paper will detail how to customize foundation models based on the pre-trained modality encoder, backbone calculator, and modality decoder. Through this approach, we can effectively leverage the knowledge of pre-trained models while reducing the computational resources demand, making the model more suitable for specific tasks and environments.
3.2.1 Transfer learning based on pre-trained modality encoders
@@
General-purpose foundation models often adapt to the distribution characteristics of a vast dataset during training, internalizing ample domain knowledge in their model parameters. They are thus well-suited as feature extraction modules for customized foundation models. By utilizing the pre-trained model’s pre-existing feature extraction module as the modality encoder for domain data, downstream task modules can be seamlessly integrated after this module to fulfill task requirements. The modality encoder encapsulates knowledge about crucial features of the data. There are two main approaches to transfer learning for obtaining modality encoders:
@@
  • Transfer within the same modality across different data domains: This approach involves transferring the knowledge of the modality encoder trained on the source data domain to the target data domain. It typically entails aligning the feature distribution of the source and target domain data. Fine-tuning the pre-trained modality encoder from the source domain on the target domain enables it to adapt to the characteristics of the new data. Specifically, this can be achieved by adjusting or adding layers to the encoder. Additionally, domain adaptation techniques such as domain adversarial training or domain-invariant feature extraction techniques can be employed to reduce the distribution discrepancy between the source and target domains.
@@
  • Transfer across different modalities: When it comes to multi-modality foundation models, there are often cases that lacking of corresponding pre-trained encoders for certain modalities. In such situations, a cross-modality transfer strategy can be utilized, adapting encoders of other modalities to process new data modalities. For instance, the authors of ImageBind treat depth and thermal imaging data as a form of single-channel images, thereby using image encoders to extract features for thermal data. Initializing the model with weights pre-trained on image datasets can lead to faster convergence compared to random initialization and can enhance generalization to some extent.
3.2.2 Transfer learning based on pre-trained backbone calculator
@@
For multi-modality foundation models, the backbone calculator is the core computational component responsible for processing encoded feature vectors and performing tasks such as classification and generation. The backbone calculator of customized foundation models can be transferred from pre-trained models to leverage the complex feature processing and task execution capabilities learned from large-scale datasets. This approach avoids training the backbone calculator from scratch but still requires constructing appropriate previous modules to encode the data into feature vectors that can be processed by the backbone calculator. For example, NExT-GPT [26] converts raw data from various modalities into language modality feature vectors before feeding them into a pre-trained LLM, allowing the model to process inputted tokens according to task requirements. There are two main approaches to transferring pre-trained backbone calculators:
@@
  • Transfer of a single pre-trained backbone calculator: Utilizing a general-purpose foundation model (e.g., LLaMA) as the backbone calculator to process data from the central modality. When transferring a pre-trained backbone calculator to a specific domain application, it typically requires fine-tuning to adapt to specific data characteristics and task requirements of that domain. This step can be performed on a limited domain dataset, fine-tuning the model parameters to optimize its processing capabilities for the domain-specific data.
  • Modular combination of multiple pre-trained backbone calculators: Modular combination is a flexible design method in deep learning architectures that allow integrating multiple specialized pre-trained models into a unified framework based on task requirements. The Mixture of Experts (MoE) model [88] can serve as an effective mechanism to further optimize this modular combination. The MoE model introduces multiple expert networks and uses a gating mechanism and mixing strategy to dynamically select and combine outputs of these experts, enabling specialized processing for different tasks or data subsets.
@@
    The primary function of the gating mechanism is to determine how input data should be distributed among different experts. It generates a weight or score for each expert based on the characteristics of the input data, reflecting each expert’s capability or suitability for processing the current input. The output of the gating mechanism is typically used to guide the mixing strategy, indicating the importance of each expert for the current input. The mixing strategy then combines the outputs of multiple experts according to specific rules to generate the final model output. This strategy can be simple, such as averaging or weighted averaging, or more complex, involving probability distributions of model outputs or other advanced methods.
@@
    For example, for complex tasks requiring both image recognition and language understanding, one expert network might excel at identifying edges in images, while another expert network might be adept at understanding semantic relationships in natural language. The gating mechanism of the MoE model can automatically adjust the participation level of each expert network based on the characteristics of the input data and the task requirement, which allows the model to flexibly invoke the most appropriate expert network when dealing with mixed visual and language inputs, achieving optimal performance. Moreover, MoE models are highly scalable. They can adapt to new task requirements or data types by adding new expert networks and updating the gating mechanism, making them suitable for constructing flexible customized foundation models.
3.2.3 Transfer learning based on pre-trained modality decoders
@@
Modality decoders play a crucial role in multi-modality foundation pre-trained models. They are responsible for converting processed feature vectors back into the form of the original data. In generative tasks, such as converting text to images or audio to text, modality decoders need not only to accurately decode the feature vectors to reconstruct understandable original data but also to exhibit a certain level of creativity. Some pre-trained modality decoders can also understand and process multi-modality feature inputs. For instance, CoDi-2 can utilize both text and audio as conditions to control image generation. By transferring such pre-trained decoders, there is no need to train complex decoder structures from scratch, allowing them to be directly applied to image generation tasks.
Here are methods to effectively utilize pre-trained modality decoders for transfer learning:
@@
    Fine-tuning pre-trained modality decoders: Similar to modality encoders, modality decoders can be fine-tuned on a specific task’s dataset to adapt to new task requirements. This process usually involves adjusting the last few layers of the decoder or adding new layers to better capture the data characteristics of the specific domain.
@@
    Transferring cross-modality generative modality decoders: In cross-modality generation tasks, pre-trained modality decoders can be directly used to generate data in the target modality. The conditional information is first encoded into feature vectors through a conditional encoder, and then combined with feature vectors of the original data to achieve conditional generation. The prerequisite for this functionality is ensuring that input feature vectors can be correctly understood by the decoder, which may involve adjustments to the backbone calculator and output projector.
@@ 3.3 Construction of the foundation model without pre-trained modules
When it is not possible to construct a foundation model through transferring from pre-trained models, it becomes necessary to design and train the corresponding modules. We will first provide a general analysis about the architectures of single-modality and multi-modality foundation models, as the foundation for constructing each components.
@@
Single-modality foundation models consist of three core modules: the modality encoder, the backbone calculator, and the modality decoder. For example, in LLaMA-2 [11], the modality encoder and decoder are specifically designed for the language modality, utilizing the byte pair encoding (BPE) algorithm for encoding and decoding functions. And the backbone calculator is a massive autoregressive Transformer model. By this way, LLaMA-2 achieves a complete processing workflow of ”input raw text – input text feature vectors – output text feature vectors – output raw text”. Additionally, Bai et al. [13] introduces the concept of visual sentences, and proposes a large visual model (LVM) that can autoregressively generate the required image based on visual sentences. It realizes in-context learning within the pure image modality, which enables the model to directly infer tasks from image modality prompts and generate corresponding results. This not only explores the potential of pure visual input but also provides a new perspective for constructing domain-specific foundation models — the central modality does not have to be limited to language but any modality widely used in a specific domain.
@@
Multi-modality foundation models require the additional input projectors and output projectors to achieve modality alignment. For instance, CoDi-2 [21] first utilizes multiple modality encoders proposed in ImageBind [25] to process input data, aligning all corresponding modalities to the image modality. Then, the feature vectors of the image modality are transformed into the feature space of the language modality through a multi-layer perceptron (MLP). In detail, it uses the pre-trained autoregressive Transformer of the LLM LLaMA-2-7b-chat-hf as the backbone calculator’s foundation. The image and audio features processed by the backbone calculator are converted back to the image domain via two MLPs. And they are then used as control vectors input of a Diffusion-based generative model to obtain the final image and text results. The training losses include text generation loss, modality conversion loss, and data generation loss. So that the multi-modality feature processing capability of the backbone calculator and the modality conversion capability of the two MLPs can be trained at the same time by an end-to-end way. The model’s modality alignment is reflected in two terms. On one hand, the model aligns the feature vectors of multiple modalities to the image modality through the pre-trained modality encoders of ImageBind. On the other hand, it also converts between image feature vectors and text feature vectors via the MLP.
@@
In summary, constructing a foundation model begins with determining the data modalities and selecting the central modality. Next, the modality encoder and input projector are implemented to convert raw data from different modalities into central modality feature vectors, which will then be processed by the backbone calculator. Following this, output projector and modality decoder are designed to convert the feature vectors from the backbone calculator back into the original data forms of each modality. Once the model structure is constructed, the training process can begin. In the following, we will provide a detailed introduction to the implementation principles and construction methods of each module.
3.3.1 Constructing modality encoders
@@
Constructing a modality encoder means designing a neural network capable of extracting feature vectors from data. The general steps for building a modality encoder are as follows:
@@
    Preprocessing into suitable data structures: Choose an appropriate data structure based on the characteristics of the data modality for subsequent model usage. For example, in audio processing, a common approach is to convert time-domain signals into spectrograms and then use neural network designed for images to extract features. And another method is to view audio as sequential vectors and use neural networks for sequence like time series or neural language to process it. When selecting the target data structure, researchers need to balance task requirements and processing difficulty, ensuring that the data structure both represents domain knowledge sufficiently and is suitable for downstream model processing. Additionally, since domain-specific foundation models need to be functionally versatile, the compatibility of various task inputs should be considered when choosing the target data structure.
@@
    Training the modality encoder: Pre-train the modality encoder using a dataset with sufficient quantity and variety of samples to learn the general features and distributions of the modality data. Pre-training is the process of infusing the model with knowledge. If the dataset size or diversity is inadequate, the model might not learn a complete representation of the modality data. One method to train a modality encoder is to combine the modality encoder and modality decoder into an autoencoder and perform unsupervised training by minimizing reconstruction error. Another training method is designing and training a model for a specific task by supervised training, and then transferring the upstream part of the model as the modality encoder. However, this method cannot get a compatible modality decoder, which might affect the design and functionality of subsequent modules. Therefore, when designing the modality encoder, it is essential to consider the consistency of the entire foundation model architecture.
3.3.2 Constructing input projectors

### v2

@@ 3.1.2 Domain-specific enhancement based on fine-tuning
      where $\Delta W$ is the Kronecker product of two low-rank matrices $A$ and $B$:
      $\Delta W=A\otimes B\text{.}$ (15)
      In this formula, two low-rank matrices, $A$ and $B$, are used by the algorithm to compute a large matrix through the Kronecker product (denoted as $\otimes$), which is then updated during the fine-tuning process. LoKr is particularly suitable for tasks that require increasing the model’s dimensions to capture more complex relationships while maintaining similar parameter efficiency to LoHa. However, LoKr may require more complex mathematical operations to handle the Kronecker product, and in some cases, its computational cost may be higher than that of LoHa.
@@
    Full fine-tuning: Full fine-tuning is not constrained by pre-training tasks or data distributions, making it flexible to adapt to various downstream tasks. It allows the model to be directly optimized end-to-end on the data of the final task without the need for additional adapter modules. However, because it requires updating all parameters in the model, full fine-tuning demands significant computational resources and longer training times. Moreover, FMs have a vast number of parameters, and insufficient fine-tuning data may lead to overfitting. Additionally, the intermediate variables generated during full fine-tuning consume a large amount of GPU memory. Therefore, researchers have proposed many parameter-efficient fine-tuning methods mentioned earlier, which can reduce resource consumption and training time while maintaining performance [110].
@@ 3.2 Customization of the foundation model based on pre-trained modules
FMs may contain millions or even billions of parameters. Through domain-specific customization, it’s possible to reduce parts of the model that need training, thereby significantly lowering training costs. This approach is known as the customization of the FM based on pre-trained modules. Such customization means utilizing the knowledge embedded in model parameters during the pre-training process when constructing a new model. As mentioned earlier, in the architecture of FMs, there are typically five main modules: MEs, IPs, BCs, OPs and MDs. Among them, the ME, BC, and MD carry a large amount of knowledge as they are directly involved in data encoding, processing, and decoding. In contrast, the IP and OP themselves carry less model knowledge. In some cases, it is unnecessary to explicitly train these modules. Therefore, when customizing FMs, the IP and OP are often excluded.
@@
Next, this paper will detail how to customize FMs based on the pre-trained ME, BC, and MD. Through this approach, we can effectively leverage pre-trained models’ knowledge while reducing the demand for computational resources, making the model more suitable for specific tasks and environments.
3.2.1 Customization of modality encoder
@@
General-purpose FMs often adapt to the distribution characteristics of a vast dataset during training, internalizing ample domain knowledge in their model parameters. They are thus well-suited as feature extraction modules for customized FMs. By utilizing the pre-trained model’s pre-existing feature extraction module as the ME for domain data, downstream task modules can be seamlessly integrated after this module to fulfill task requirements. The ME encapsulates knowledge about crucial features of the data. There are two main approaches to customizing MEs:
Figure 10: An example of using LLM as the pre-trained BC to adopt to the autonomous driving area. In this example, multi-view camera data and LiDAR data are sent into a vision encoder aligned with LLM’s feature space. By incorporating the user’s instructions and environmental information, the LLM can generate real-time control signals for the car. The data used in the example is from [122].
@@
  • Customization within the same modality across different data domains: This approach involves transferring the knowledge of the ME trained on the source data domain to the target data domain. It typically entails aligning the feature distribution of the source and target domain data. Fine-tuning the pre-trained ME from the source domain on the target domain enables it to adapt to the characteristics of the new data. Specifically, this can be achieved by adjusting or adding layers to the encoder. Additionally, domain adaptation techniques such as domain adversarial training or domain-invariant feature extraction techniques can reduce the distribution discrepancy between the source and target domains.
@@
  • Customization across modalities: When it comes to multi-modality FMs, there are often cases that lack corresponding pre-trained encoders for certain modalities. In such situations, a cross-modality customization strategy can be utilized, adapting encoders to process new data modalities similar to the original modality. Although the modality is different, the source and target modality with the same data structure, such as natural images and thermal images, can be simultaneously fed into one ME. Such customization involves fine-tuning a pre-trained ME using task-specific datasets. For instance, ImageBind treats depth and thermal imaging data as single-channel images, using image encoders to extract features for thermal data. Initializing the model with weights pre-trained on image datasets can lead to faster convergence than random initialization and enhance generalization. This strategy is particularly useful in applications like autonomous driving, where vehicles need to process diverse data modalities, including visual, depth, thermal, and LiDAR (Light Detection And Ranging) data. To more directly understand this process, we present another example in Figure 10. In this example, natural images captured by multi-view cameras are fed into the pre-trained vision encoder with LiDAR data.
    By adapting image encoders to handle thermal imaging, the model can better understand and process thermal data, improving the overall performance of the autonomous driving system. This cross-modal customization increases the flexibility of multi-modal systems and enhances their adaptability and robustness, providing a powerful tool for solving complex problems.
3.2.2 Customization of backbone calculator
@@
For multi-modality FMs, the BC is the core computational component for processing encoded feature vectors and performing tasks such as classification and generation. The BC of customized FMs can be customized from pre-trained models to leverage the complex feature processing and task execution capabilities learned from large-scale datasets. This approach avoids training the BC from scratch but requires constructing appropriate previous modules to encode the data into feature vectors that the BC can process. For example, NExT-GPT [26] converts raw data from various modalities into language modality feature vectors before feeding them into a pre-trained LLM, allowing the model to process inputted tokens according to task requirements. There are two main approaches to customize pre-trained BCs:
@@
  • Customization from a single pre-trained BC: Utilizing a general-purpose FM (e.g., LLaMA) as the BC to process data from the central modality. Customizing a pre-trained BC to a specific domain application typically requires fine-tuning to adapt to domain-specific data characteristics and task requirements. This step can be performed on a limited domain dataset, fine-tuning the model parameters to optimize its processing capabilities for the domain-specific data.
  • Modular combination of multiple pre-trained BCs: Modular combination is a flexible design method in deep learning architectures that allows integrating multiple specialized pre-trained models into a unified framework based on task requirements. The Mixture of Experts (MoE) model [123] can serve as an effective mechanism to further optimize this modular combination. The MoE model introduces multiple expert networks and uses a gating mechanism and mixing strategy to dynamically select and combine the outputs of these experts, enabling specialized processing for different tasks or data subsets. Recent advancements in large-scale foundation models have adopted sparse MoE architectures to achieve a favorable balance between model capacity and computational efficiency. For instance, DeepSeek-V3 employs a large-scale MoE setup with a total of 671 billion parameters, yet activates only 37 billion parameters per token using a top-2 expert routing strategy [124]. It introduces an efficient expert-balancing mechanism without auxiliary loss, and leverages a multi-token prediction objective for improved generalization. Similarly, Mixtral-8x7B [125] activates only two out of eight expert subnetworks (each a 7B model) per forward pass, enabling performance comparable to much larger dense models while significantly reducing inference costs. These examples demonstrate how sparse MoE designs have become core architectural elements in state-of-the-art foundation models, offering scalable specialization with manageable computational overhead.
@@
    The primary function of the gating mechanism is to determine how input data should be distributed among experts. It generates a weight or score for each expert based on the characteristics of the input data, reflecting each expert’s capability or suitability for processing the current input. The output of the gating mechanism is typically used to guide the mixing strategy, indicating the importance of each expert for the current input. The mixing strategy then combines the outputs of multiple experts according to specific rules to generate the final model output. This strategy can be simple, such as averaging or weighted averaging, or more complex, involving probability distributions of model outputs or other advanced methods.
@@
    For example, for complex tasks requiring both image recognition and language understanding, one expert network might excel at identifying edges in images. In contrast, another expert network might be adept at understanding semantic relationships in natural language. The MoE model’s gating mechanism can automatically adjust each expert network’s participation level based on the input data’s characteristics and the task requirement, allowing the model to flexibly invoke the most appropriate expert network when dealing with mixed visual and language inputs, achieving optimal performance. Moreover, MoE models are highly scalable. They can adapt to new task requirements or data types by adding new expert networks and updating the gating mechanism, making them suitable for constructing flexible customized FMs.
3.2.3 Customization of modality decoder
@@
MDs play a crucial role in multi-modality foundation pre-trained models for converting processed feature vectors back into the form of the original data. In generative tasks, such as converting text to images or audio to text, MDs need to accurately decode the feature vectors to reconstruct understandable original data and exhibit a certain level of creativity. Some pre-trained MDs can also understand and process multi-modality feature inputs. For instance, CoDi-2 can utilize both text and audio as conditions to control image generation. By customizing such pre-trained decoders, there is no need to train complex decoder structures from scratch, allowing them to be directly applied to image generation tasks.
Here are methods to effectively utilize pre-trained MDs for customization:
@@
    Fine-tuning pre-trained MDs: Similar to MEs, MDs can be fine-tuned on a specific task’s dataset to adapt to new task requirements. This process usually involves adjusting the last few layers of the decoder or adding new layers to better capture the data characteristics of the specific domain.
@@
    Customizing cross-modality generative MDs: In cross-modality generation tasks, pre-trained MDs can be directly used to generate data in the target modality. The conditional information is first encoded into feature vectors through a conditional encoder and then combined with feature vectors of the original data to achieve conditional generation. The prerequisite for this functionality is ensuring that the decoder can correctly understand input feature vectors, which may involve adjustments to the BC and OP.
@@ 3.3 Construction of the foundation model without pre-trained modules
When it is not possible to construct an FM through transferring from pre-trained models, it becomes necessary to design and train the corresponding modules. We will first provide a general analysis of the architectures of single-modality and multi-modality FMs as the foundation for constructing each component.
@@
Single-modality FMs consist of three core modules: the ME, the BC, and the MD. For example, in LLaMA-2 [11], the ME and decoder are specifically designed for the language modality, utilizing the byte pair encoding (BPE) algorithm for encoding and decoding functions. And the BC is a massive autoregressive Transformer model. In this way, LLaMA-2 achieves a complete processing workflow of ”input raw text – input text feature vectors – output text feature vectors – output raw text.” Additionally, Bai et al. [13] introduce the concept of visual sentences and propose a large visual model (LVM) that can autoregressively generate the required image based on visual sentences. It realizes in-context learning within the pure image modality, which enables the model to infer tasks directly from image modality prompts and generate corresponding results. This not only explores the potential of pure visual input but also provides a new perspective for constructing domain-specific FMs — the central modality does not have to be limited to language but any modality widely used in a specific domain.
@@
Multi-modality FMs require the additional IPs and OPs to achieve modality alignment. For instance, CoDi-2 [21] first utilizes multiple MEs proposed in ImageBind [25] to process input data, aligning all corresponding modalities to the image modality. Then, the feature vectors of the image modality are transformed into the feature space of the language modality through a MLP. In detail, it uses the pre-trained autoregressive Transformer of the LLM LLaMA-2-7b-chat-hf as the BC’s foundation. The image and audio features processed by the BC are converted back to the image domain via two MLPs. They are then used as control vector input of a Diffusion-based generative model to obtain the final image and text results. The training losses include text generation, modality conversion, and data generation loss. So that the multi-modality feature processing capability of the BC and the modality conversion capability of the two MLPs can be trained simultaneously in an end-to-end way. The model’s modality alignment is reflected in two terms. On the one hand, the model aligns the feature vectors of multiple modalities to the imaging modality through the pre-trained MEs of ImageBind. On the other hand, it also converts between image feature vectors and text feature vectors via the MLP.
@@
In summary, constructing an FM begins with determining the data modalities and selecting the central modality. Next, the ME and IP are implemented to convert raw data from different modalities into central modality feature vectors, which the BC will then process. Following this, OP and MD are designed to convert the feature vectors from the BC back into the original data forms of each modality. Once the model structure is constructed, the training process can begin. In the following, we will provide a detailed introduction to each module’s implementation principles and construction methods.
3.3.1 Constructing modality encoder
@@
Constructing a ME means designing a neural network capable of extracting feature vectors from data. The general steps for building a ME are as follows:
@@
    Preprocessing into suitable data structures: Choose an appropriate data structure based on the characteristics of the data modality for subsequent model usage. For example, in audio processing, a common approach is to convert time-domain signals into spectrograms and then use a neural network designed for images to extract features. Another method is to view audio as sequential vectors and use neural networks for sequences, like time series or neural language, to process it. When selecting the target data structure, researchers need to balance task requirements and processing difficulty, ensuring that the data structure represents domain knowledge sufficiently and is suitable for downstream model processing. Additionally, since domain-specific FMs need to be functionally versatile, the compatibility of various task inputs should be considered when choosing the target data structure.
@@
    Training the ME: Pre-train the ME using a dataset with sufficient quantity and variety of samples to learn the general features and distributions of the modality data. Pre-training is the process of infusing the model with knowledge. If the dataset size or diversity is inadequate, the model might not learn a complete representation of the modality data. One method to train an ME is to combine the ME and MD into an autoencoder and perform unsupervised training by minimizing reconstruction error. Another training method is designing and training a model for a specific task by supervised training, then transferring the model’s upstream part as the ME. However, this method cannot get a compatible MD, which might affect the design and functionality of subsequent modules. Therefore, when designing the ME, it is essential to consider the consistency of the entire FM architecture.
3.3.2 Constructing input projector